# On relative tractor bundles

### v1

@@ 2.2. The group level and invariant filtrations
It is usual in the theory of parabolic geometries, that gradings have to be viewed as auxiliary objects, since they are not invariant under the actions of parabolic subalgebras. To obtain objects that are invariant, one has to pass to associated filtrations. To do this, we first recall from [Rel-BGG1] how to appropriately choose groups in the situation of nested parabolics. Given $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$ as above, we first fix a group $G$ with Lie algebra $\mathfrak{g}$ and choose a parabolic subgroup $P\subset G$ corresponding to $\mathfrak{p}$. It is well known that this means that $P$ lies between the normalizer $N_{G}(\mathfrak{p})$ of $\mathfrak{p}$ in $G$ and its connected component of the identity. Finally, we choose a subgroup $Q\subset P$ corresponding to the Lie subalgebra $\mathfrak{q}$, which means that it lies between $N_{P}(\mathfrak{q})$ and its connected component of the identity. Since $P$ and $Q$ are parabolic subgroups, it is well known that the exponential map restricts to diffeomorphisms from $\mathfrak{p}_{+}$ onto a closed normal subgroup $P_{+}\subset P$ and from $\mathfrak{q}_{+}$ onto a closed normal subgroup $Q_{+}\subset Q$. From Proposition 2, we know that $\mathfrak{p}_{+}\subset\mathfrak{q}_{+}$ so $P_{+}\subset Q_{+}\subset Q$.
@@
Parabolic geometries of type $(G,Q)$ come with principal $Q$-bundles, so via the construction of associated vector bundles, any representation of $Q$ determines a natural vector bundle on each manifold endowed with such a geometry. Observe the by construction the subgroup $Q_{+}\subset Q$ is nilpotent and it is well known that the quotient $Q/Q_{+}$ is reductive. Hence representations of $Q$ are not easy to understand in general. Special classes of natural bundles are obtained via considering a restricted class of representations of $Q$, which often are easier to understand. Classical examples of this situation (which are available for any parabolic subgroup) are completely reducible bundles that correspond to direct sums of irreducible representations of $Q$ (on which $Q_{+}$ automatically acts trivially) and tractor bundles that correspond to restrictions to $Q$ of representations of the group $G$. So in both these cases the relevant representation theory is well understood.
@@
In the more restrictive situation of two nested parabolics $Q\subset P\subset G$, two more classes of natural bundles were introduced in Definition 2.1 of [Rel-BGG2]. These are relative natural bundles which correspond to representations of $Q$ on which $P_{+}$ acts trivially and relative tractor bundles which correspond to restrictions to $Q$ of representations of $P$ on which $P_{+}$ acts trivially. So in particular any completely reducible representation of $P$ gives rise to a relative tractor bundle on geometries of type $(G,Q)$. There is a simple family of appropriate representations of $P$, which are just the filtration components of the filtration of $\mathfrak{g}$ determined by $\mathfrak{p}$.
Definition 3.
@@
Consider a real or complex semisimple Lie algebra $\mathfrak{g}$ endowed with the bigrading $\mathfrak{g}=\oplus_{i^{\prime},i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ determined by two nested standard parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$. Then for each $i^{\prime}$, we define $\mathfrak{g}^{(i^{\prime},*)}:=\oplus_{j^{\prime}\geq i^{\prime},j^{\prime\prime}}\mathfrak{g}_{(j^{\prime},j^{\prime\prime})}$.
@@
Proposition 4.
Consider a real or complex semisimple Lie algebra $\mathfrak{g}$ endowed with the bigrading $\mathfrak{g}=\oplus_{i^{\prime},i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ determined by two nested standard parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$. Then for each integer $i^{\prime}$ the subspace $\mathfrak{g}^{(i^{\prime},*)}\subset\mathfrak{g}$ is invariant for the restriction of the adjoint action of $G$ to $P$. Moreover, $[\mathfrak{p}_{+},\mathfrak{g}^{(i^{\prime},*)}]\subset\mathfrak{g}^{(i^{\prime}+1,*)}$, so the subgroup $P_{+}$ acts trivially on the quotient $\mathbb{V}_{i^{\prime}}:=\mathfrak{g}^{(i^{\prime},*)}/\mathfrak{g}^{(i^{\prime}+1,*)}$.
For each $i^{\prime\prime}$ such that $\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}\neq 0$, the subspace $\oplus_{j^{\prime}\geq i^{\prime},j^{\prime\prime}\geq i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},j^{\prime})}$ descend to a subspace $\mathbb{V}^{i^{\prime\prime}}_{i^{\prime}}\subset\mathbb{V}_{i^{\prime}}$. These subspace define a $Q$-invariant decreasing filtration of $\mathbb{V}_{i^{\prime}}$ in the sense that for $i^{\prime\prime}\leq j^{\prime\prime}$ we get $\mathbb{V}^{i^{\prime\prime}}_{i^{\prime}}\subset\mathbb{V}^{j^{\prime\prime}}_{i^{\prime}}$.
@@
The grading of $\mathfrak{g}$ induced by $\mathfrak{p}$ induces a filtration, whose components by construction are exactly the space $\mathfrak{g}^{(i^{\prime},*)}$. It is well known this components can be obtained algebraically from $\mathfrak{p}$ as the nilradical, the elements of its derived series and annihilators under the Killing form, compare with Corollary 3.2.1 of [book]. Hence any element of $N_{G}(\mathfrak{p})$ normalizes each of the filtration components which shows that each $\mathfrak{g}^{(i^{\prime},*)}$ is $P$-invariant. In particular, $\mathfrak{p}_{+}=\mathfrak{g}^{(1,*)}$ and the second claim follows from Proposition 2. Applying this argument to $Q$, we conclude that each of the spaces $\oplus_{j^{\prime}\geq i^{\prime},j^{\prime\prime}\geq i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},j^{\prime})}$ is $Q$-invariant and from this the last claim readily follows. ∎
2.3. Some basic relative tractor bundles
@@
We can now obtain the distinguished relative tractor bundles by passing to associated bundles. Recall that, given a parabolic geometry $(p:\mathcal{G}\to M,\omega)$, the associated bundle induced by the restriction of the adjoint representation of $G$ to $Q$ is the adjoint tractor bundle $\mathcal{A}M=\mathcal{G}\times_{Q}\mathfrak{g}$. By Proposition 4, we see that for each $i^{\prime}$, we get a natural subbundle corresponding to the subspace $\mathfrak{g}^{(i^{\prime},*)}\subset\mathfrak{g}$ which is $P$-invariant and hence $Q$ invariant. We denote this subbundle by $\mathcal{A}^{(i^{\prime},*)}M\subset\mathcal{A}M$. Of course we can then form the quotient $\mathcal{V}_{i^{\prime}}M:=\mathcal{A}^{(i^{\prime},*)}M/\mathcal{A}^{(i^{\prime}+1,*)}M$ which is isomorphic to $\mathcal{G}\times_{Q}(\mathfrak{g}^{(i^{\prime},*)}/\mathfrak{g}^{(i^{\prime}+1,*)})=\mathcal{G}\times_{Q}\mathbb{V}_{i^{\prime}}$. Hence by Proposition 4, each $\mathbb{V}_{i^{\prime}}$ is a relative tractor bundle. Moreover, the $Q$-invariant subspaces $\mathbb{V}_{i^{\prime}}^{i^{\prime\prime}}\subset\mathbb{V}_{i^{\prime}}$ give rise to a filtration of $\mathcal{V}_{i^{\prime}}M$ smooth subbundles $\mathcal{V}_{i^{\prime}}^{i^{\prime\prime}}M$, each of which is a relative natural bundle.
@@
In what follows, we will mainly be interested in the case that $i^{\prime}<0$. The relevance of this condition is that by construction $\mathfrak{g}^{(0,*)}=\mathfrak{p}\supset\mathfrak{q}$. It is well known that, via the Cartan connection $\omega$, the associated bundle $\mathcal{G}\times_{Q}\mathfrak{g}/\mathfrak{q}$ is isomorphic to $TM$. Now the subspace $\mathfrak{p}\subset\mathfrak{g}$ is $P$-invariant and hence $Q$-invariant, so $\mathfrak{p}/\mathfrak{q}\subset\mathfrak{g}/\mathfrak{q}$ is a $Q$-invariant subspace, too. Hence it defines a natural subbundle in $TM$, which is called the relative tangent bundle $T_{\rho}M$ in [Rel-BGG2]. For $i^{\prime}<0$, we get in the same way a $Q$-invariant subspace $\mathfrak{g}^{(i^{\prime},*)}/\mathfrak{q}$ which gives rise to a smooth subbundle of $TM$ that contains $T_{\rho}M$. We denote this bundle by $T^{i^{\prime}}_{P}M$ to distinguish it form the filtration component coming from the initial parabolic geometry. By construction, we then can naturally identify $\mathcal{V}_{-1}M$ with $T^{-1}_{P}M/T_{\rho}M$ and, for $i^{\prime}<-1$, $\mathcal{V}_{i^{\prime}}M=T^{i^{\prime}}_{P}M/T^{i^{\prime}+1}_{P}M$, so we always obtain natural subquotients of the tangent bundle. This will be crucial for the later developments. In what follows it will be convenient to define $T^{0}_{P}M:=T_{\rho}M$, which in particular allows us to uniformly write $\mathcal{V}_{i^{\prime}}M=T^{i^{\prime}}_{P}M/T^{i^{\prime}+1}_{P}M$ for any $i^{\prime}\leq-1$.
@@
A particularly important special case is that the relative tangent bundle $T_{\rho}M\subset TM$ is an involutive distribution on $M$. In this case, we can find local leaf spaces for this distribution. More explicitly, for any $x\in M$ there is an open neighborhood $U\subset M$ of $x$ and a surjective submersion $\psi:U\to N$ onto some smooth manifold $N$ such that, for any $y\in U$, $\ker(T_{y}\psi)$ coincides with the fiber of $T_{\rho}M$ at $y$. In particular, this implies that $T\psi$ induces an isomorphism $(TM/T_{\rho}M)|_{U}\cong\psi^{*}TN$. Characterizing involutivity of $T_{\rho}M$ and the question of which part of the given geometric structure on $M$ descends to local leaf spaces is a key part of the theory of twistor spaces for parabolic geometries that was originally developed in [twistor]. Recall that the curvature of the Cartan connection $\omega$ can be interpreted as a two-form $\kappa\in\Omega^{2}(M,\mathcal{A}M)$ and since $TM$ naturally is a quotient of $\mathcal{A}M$, there is a projection $\tau\in\Omega^{2}(M,TM)$ of $\kappa$, called the torsion of the Cartan connection $\omega$. By Proposition 4.6 of [Rel-BGG2], involutivity of $T_{\rho}M$ is equivalent to the fact that $\tau$ maps $T_{\rho}M\times T_{\rho}M$ to $T_{\rho}M$.
@@
There is a strengthening of this condition that plays an important role in the theory of relative BGG sequences, namely that $\kappa$ vanishes on $T_{\rho}M\times T_{\rho}M$. (In [Rel-BGG2], this is phrased as vanishing of the relative curvature.) By Theorem 4.11 of [Rel-BGG2], this implies that the relative BGG sequence determined by any relative tractor bundle is a complex and a fine resolution of a sheaf on $M$. This sheaf can be identified with the sheaf of those local sections of the relative tractor bundle which are parallel for the relative tractor connection, which will be discussed in §3.1 below. It is shown in Section 4.7 of [Rel-BGG2] that this implies that locally the resolved sheaf can be written as a pullback of a sheaf on a local leaf space. In the special case of a so-called correspondence space (which means that the geometry descends to the leaf space), that sheaf is identified explicitly in Theorem 4.11 of [Rel-BGG2]. In general, this sheaf is not identified in [Rel-BGG2], we’ll present a solution to this question under some assumptions below.
3. Relative tractor connections and resolved sheaves
3.1. Restrictions of the adjoint tractor connection

### v2

@@ 2.2 The group level and invariant filtrations
It is usual in the theory of parabolic geometries, that gradings have to be viewed as auxiliary objects, since they are not invariant under the actions of parabolic subalgebras. To obtain objects that are invariant, one has to pass to associated filtrations. To do this, we first recall from [9] how to appropriately choose groups in the situation of nested parabolics. Given $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$ as above, we first fix a group $G$ with Lie algebra $\mathfrak{g}$ and choose a parabolic subgroup $P\subset G$ corresponding to $\mathfrak{p}$. It is well known that this means that $P$ lies between the normalizer $N_{G}(\mathfrak{p})$ of $\mathfrak{p}$ in $G$ and its connected component of the identity. Finally, we choose a subgroup $Q\subset P$ corresponding to the Lie subalgebra $\mathfrak{q}$, which means that it lies between $N_{P}(\mathfrak{q})$ and its connected component of the identity. Since $P$ and $Q$ are parabolic subgroups, it is well known that the exponential map restricts to diffeomorphisms from $\mathfrak{p}_{+}$ onto a closed normal subgroup $P_{+}\subset P$ and from $\mathfrak{q}_{+}$ onto a closed normal subgroup $Q_{+}\subset Q$. From Proposition 2.2, we know that $\mathfrak{p}_{+}\subset\mathfrak{q}_{+}$ so $P_{+}\subset Q_{+}\subset Q$.
@@
Parabolic geometries of type $(G,Q)$ come with principal $Q$-bundles, so via the construction of associated vector bundles, any representation of $Q$ determines a natural vector bundle on each manifold endowed with such a geometry. Observe that by construction the subgroup $Q_{+}\subset Q$ is nilpotent and it is well known that the quotient $Q/Q_{+}$ is reductive. Hence, representations of $Q$ are not easy to understand in general. Special classes of natural bundles are obtained via considering a restricted class of representations of $Q$, which often are easier to understand. Classical examples of this situation (which are available for any parabolic subgroup) are completely reducible bundles that correspond to direct sums of irreducible representations of $Q$ (on which $Q_{+}$ automatically acts trivially) and tractor bundles that correspond to restrictions to $Q$ of representations of the group $G$. So in both these cases the relevant representation theory is well understood.
@@
In the more restrictive situation of two nested parabolics $Q\subset P\subset G$, two more classes of natural bundles were introduced in [10, Definition 2.1]. These are relative natural bundles which correspond to representations of $Q$ on which $P_{+}$ acts trivially and relative tractor bundles which correspond to restrictions to $Q$ of representations of $P$ on which $P_{+}$ acts trivially. So in particular any completely reducible representation of $P$ gives rise to a relative tractor bundle on geometries of type $(G,Q)$. There is a simple family of appropriate representations of $P$, which are just the filtration components of the filtration of $\mathfrak{g}$ determined by $\mathfrak{p}$.
Definition 2.3.
@@
Consider a real or complex semisimple Lie algebra $\mathfrak{g}$ endowed with the bigrading $\mathfrak{g}=\bigoplus_{i^{\prime},i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ determined by two nested standard parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$. Then for each $i^{\prime}$, we define $\mathfrak{g}^{(i^{\prime},*)}:=\bigoplus_{j^{\prime}\geq i^{\prime},j^{\prime\prime}}\mathfrak{g}_{(j^{\prime},j^{\prime\prime})}$.
@@
Proposition 2.4.
Consider a real or complex semisimple Lie algebra $\mathfrak{g}$ endowed with the bigrading $\mathfrak{g}=\bigoplus_{i^{\prime},i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ determined by two nested standard parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$. Then for each integer $i^{\prime}$ the subspace $\mathfrak{g}^{(i^{\prime},*)}\subset\mathfrak{g}$ is invariant for the restriction of the adjoint action of $G$ to $P$. Moreover, $\bigl{[}\mathfrak{p}_{+},\mathfrak{g}^{(i^{\prime},*)}\bigr{]}\subset\mathfrak{g}^{(i^{\prime}+1,*)}$, so the subgroup $P_{+}$ acts trivially on the quotient $\mathbb{V}_{i^{\prime}}:=\mathfrak{g}^{(i^{\prime},*)}/\mathfrak{g}^{(i^{\prime}+1,*)}$. For each $i^{\prime\prime}$ such that $\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}\neq 0$, let $\mathbb{V}^{i^{\prime\prime}}_{i^{\prime}}\subset\mathbb{V}_{i^{\prime}}$ be the image of the subspace $\bigoplus_{j^{\prime}\geq i^{\prime},j^{\prime\prime}\geq i^{\prime\prime}}\mathfrak{g}_{(j^{\prime},j^{\prime\prime})}\subset\mathfrak{g}^{(i^{\prime},*)}$ under the quotient projection. Then each of the spaces $\mathbb{V}^{i^{\prime\prime}}_{i^{\prime}}$ is $Q$-invariant and they define a decreasing filtration of $\mathbb{V}_{i^{\prime}}$ in the sense that for $i^{\prime\prime}\leq\ell^{\prime\prime}$ we get ${\mathbb{V}^{i^{\prime\prime}}_{i^{\prime}}\supset\mathbb{V}^{\ell^{\prime\prime}}_{i^{\prime}}}$.
@@
The grading of $\mathfrak{g}$ induced by $\mathfrak{p}$ induces a filtration, whose components by construction are exactly the space $\mathfrak{g}^{(i^{\prime},*)}$. It is well known this components can be obtained algebraically from $\mathfrak{p}$ as the nilradical, the elements of its derived series and annihilators under the Killing form, compare with [7, Corollary 3.2.1]. Hence, any element of $N_{G}(\mathfrak{p})$ normalizes each of the filtration components which shows that each $\mathfrak{g}^{(i^{\prime},*)}$ is $P$-invariant. In particular, $\mathfrak{p}_{+}=\mathfrak{g}^{(1,*)}$ and the second claim follows from Proposition 2.2. Applying this argument to $Q$, we conclude that each of the spaces $\bigoplus_{j^{\prime}\geq i^{\prime},j^{\prime\prime}\geq i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},j^{\prime})}$ is $Q$-invariant and from this the second part follows readily. ∎
2.3 Some basic relative tractor bundles
@@
We can now obtain the distinguished relative tractor bundles by passing to associated bundles. Recall that, given a parabolic geometry $(p\colon\mathcal{G}\to M,\omega)$, the associated bundle induced by the restriction of the adjoint representation of $G$ to $Q$ is the adjoint tractor bundle $\mathcal{A}M=\mathcal{G}\times_{Q}\mathfrak{g}$. By Proposition 2.4, we see that for each $i^{\prime}$, we get a natural subbundle corresponding to the subspace $\mathfrak{g}^{(i^{\prime},*)}\subset\mathfrak{g}$ which is $P$-invariant and hence $Q$ invariant. We denote this subbundle by $\mathcal{A}^{(i^{\prime},*)}M\subset\mathcal{A}M$. Of course we can then form the quotient $\mathcal{V}_{i^{\prime}}M:=\mathcal{A}^{(i^{\prime},*)}M/\mathcal{A}^{(i^{\prime}+1,*)}M$ which is isomorphic to $\mathcal{G}\times_{Q}\bigl{(}\mathfrak{g}^{(i^{\prime},*)}/\mathfrak{g}^{(i^{\prime}+1,*)}\bigr{)}=\mathcal{G}\times_{Q}\mathbb{V}_{i^{\prime}}$. Hence, by Proposition 2.4, each $\mathbb{V}_{i^{\prime}}$ is a relative tractor bundle. Moreover, the $Q$-invariant subspaces $\mathbb{V}_{i^{\prime}}^{i^{\prime\prime}}\subset\mathbb{V}_{i^{\prime}}$ give rise to a filtration of $\mathcal{V}_{i^{\prime}}M$ by smooth subbundles $\mathcal{V}_{i^{\prime}}^{i^{\prime\prime}}M$, each of which is a relative natural bundle.
@@
In what follows, we will mainly be interested in the case where $i^{\prime}<0$. The relevance of this condition is that by construction $\mathfrak{g}^{(0,*)}=\mathfrak{p}\supset\mathfrak{q}$. It is well known that, via the Cartan connection $\omega$, the associated bundle $\mathcal{G}\times_{Q}\mathfrak{g}/\mathfrak{q}$ is isomorphic to $TM$. Now the subspace $\mathfrak{p}\subset\mathfrak{g}$ is $P$-invariant and hence $Q$-invariant, so $\mathfrak{p}/\mathfrak{q}\subset\mathfrak{g}/\mathfrak{q}$ is a $Q$-invariant subspace, too. Hence, it defines a natural subbundle in $TM$, which is called the relative tangent bundle $T_{\rho}M$ in [10]. For $i^{\prime}<0$, we get in the same way a $Q$-invariant subspace $\mathfrak{g}^{(i^{\prime},*)}/\mathfrak{q}$ which gives rise to a smooth subbundle of $TM$ that contains $T_{\rho}M$. We denote this bundle by $T^{i^{\prime}}_{P}M$ to distinguish it form the filtration component coming from the initial parabolic geometry. By construction, we then can naturally identify $\mathcal{V}_{-1}M$ with $T^{-1}_{P}M/T_{\rho}M$ and, for $i^{\prime}<-1$, $\mathcal{V}_{i^{\prime}}M=T^{i^{\prime}}_{P}M/T^{i^{\prime}+1}_{P}M$, so we always obtain natural subquotients of the tangent bundle. This will be crucial for the later developments. In what follows it will be convenient to define $T^{0}_{P}M:=T_{\rho}M$, which in particular allows us to uniformly write $\mathcal{V}_{i^{\prime}}M=T^{i^{\prime}}_{P}M/T^{i^{\prime}+1}_{P}M$ for any $i^{\prime}\leq-1$.
@@
A particularly important special case is that the relative tangent bundle $T_{\rho}M\subset TM$ is an involutive distribution on $M$. In this case, we can find local leaf spaces for this distribution. More explicitly, for any $x\in M$ there is an open neighborhood $U\subset M$ of $x$ and a surjective submersion $\psi\colon U\to N$ onto some smooth manifold $N$ such that, for any $y\in U$, $\ker(T_{y}\psi)$ coincides with the fiber of $T_{\rho}M$ at $y$. In particular, this implies that $T\psi$ induces an isomorphism $(TM/T_{\rho}M)|_{U}\cong\psi^{*}TN$. Characterizing involutivity of $T_{\rho}M$ and the question of which part of the given geometric structure on $M$ descends to local leaf spaces is a key part of the theory of twistor spaces for parabolic geometries that was originally developed in [5]. Recall that the curvature of the Cartan connection $\omega$ can be interpreted as a two-form $\kappa\in\Omega^{2}(M,\mathcal{A}M)$ and since $TM$ naturally is a quotient of $\mathcal{A}M$, there is a projection $\tau\in\Omega^{2}(M,TM)$ of $\kappa$, called the torsion of the Cartan connection $\omega$. By [10, Proposition 4.6], involutivity of $T_{\rho}M$ is equivalent to the fact that $\tau$ maps $T_{\rho}M\times T_{\rho}M$ to $T_{\rho}M$.
@@
There is a strengthening of this condition that plays an important role in the theory of relative BGG sequences, namely that $\kappa$ vanishes on $T_{\rho}M\times T_{\rho}M$. (In [10], this is phrased as vanishing of the relative curvature.) By [10, Theorem 4.11], this implies that the relative BGG sequence determined by any relative tractor bundle is a complex and a fine resolution of a sheaf on $M$. This sheaf can be identified with the sheaf of those local sections of the relative tractor bundle which are parallel for the relative tractor connection, which will be discussed in Section 3.1 below. It is shown in [10, Section 4.7] that this implies that locally the resolved sheaf can be written as a pullback of a sheaf on a local leaf space. In the special case of a so-called correspondence space (which means that the geometry descends to the leaf space), that sheaf is identified explicitly in [10, Theorem 4.11]. In general, this sheaf is not identified in [10], we will present a solution to this question under some assumptions below.
3 Relative tractor connections and resolved sheaves
3.1 Restrictions of the adjoint tractor connection